A tank measuring 2m times 50cm times im is 3/5 filled with water.How many litres of water must be added to fully fill the tank?
1
Expert's answer
2011-09-18T15:46:00-0400
One has to add 2/5 of total voulme of tank, which constitute:

2 * 0.5 * 1
* 2/5 = 2/5& = 0.4 m^3 = 0.4 * 1000 litres = 400 litres.